I want to make sure I am correct on this answer to the inequality. The inequality equation is
2x-5/12 + 3-x/4 is < or = to 4/3
The answer I got was x is = to or > -3/-2
(* the "/" are meant to put the problem in fraction form, like 4 over 5: 4/5)
Thank You so much!

Answer by rothauserc(4718)

2x-5/12 + 3-x/4 is < or = to 4/3
a) I assume you mean
(2x-5)/12 + (3-x)/4 is < or = to 4/3
multiply both sides of < or = by 12
2x-5 + 9-3x is < or = 16
-x +4 is < or = 16
subtract 4 from both sides of < or =
-x < or = 12
x > or = -12
b) I assume no parenthesis
2x-5/12 + 3-x/4 is < or = to 4/3
multiply both sides of < or = by 12
24x-5+36-3x is < or = 16
21x+31 is < or = 16
21x < or = -15
x < or = -5/7